Find the first quandrant area under the curve y=xe^-x.


1
Expert's answer
2021-04-13T13:30:43-0400

Let us find the first quandrant area under the curve "y=xe^{-x}". For this firstly let us sketch the graph of "y=xe^{-x}:"




It follows that the area "A" is


"A=\int_0^{+\infty}xe^{-x}dx=\lim\limits_{a\to +\infty}\int_0^{a}xe^{-x}dx="


"|u=x, dv=e^{-x}dx,du=dx, v=-e^{-x}|"


"=\lim\limits_{a\to +\infty}(-xe^{-x}|_0^a+\int_0^{a}e^{-x}dx)
=\lim\limits_{a\to +\infty}(-ae^{-a}-e^{-x}|_0^a)=
\lim\limits_{a\to +\infty}(-ae^{-a}-e^{-a}+1)=
\lim\limits_{a\to +\infty}\frac{-a-1}{e^a}+1="


| let us use the L'Hôpital's Rule |


"=\lim\limits_{a\to +\infty}\frac{-1}{e^a}+1=1"